OpenStudy (anonymous):

3.log(3x+25)=2 1.25 2.75 3.2/3 4.100 4.16^2x=124 1.0.869 2.1.150 3.1.739 4.3.477

OpenStudy (anonymous):

3.log(3x+25)=2 (3x+25)=10^2 . . . 4.16^2x=124 2x*ln(16) = ln(124) . . . does that help?